OpenStudy (anonymous):

compounded interest at 8% annually for 20 months do i divide the percent by months or vice versa? thanks in advance

OpenStudy (jim766):

A= p ( 1 + r)^t Need to know how many yrs 20 myths is

OpenStudy (jim766):

Change the percent to a decimal 8/100 Use that in the formula

OpenStudy (jim766):

.08

OpenStudy (anonymous):

8% annually and 6% quartley both 5 yrs

OpenStudy (anonymous):

can you just explain where the 100 came in to play

OpenStudy (jim766):

Above will work for annually A= p( 1 +(r/4))^4t For quarterly

OpenStudy (anonymous):

sorry still learning this getting old now at 57 to get my degree lol

OpenStudy (anonymous):

but they do say its never to late lol

OpenStudy (jim766):

Think of 1/2 = .5 To change a decimal to a%, we then mult by 100 To get 50% To change a percent to a decimal, we do just the opposite And divide by 100

OpenStudy (jim766):

You are right never too late

OpenStudy (anonymous):

that part i understand

OpenStudy (anonymous):

what type of calculator do i need?

OpenStudy (jim766):

In the formulas above the p is the principle, the amt you start out with, Just substitute the info in and solve

OpenStudy (anonymous):

its that type of formula i am still trying to learn

OpenStudy (anonymous):

sorry for being slow in understanding

OpenStudy (anonymous):

i appreciate the help

OpenStudy (jim766):

In the annual problem how much do they tell you to start with?

OpenStudy (anonymous):

500000 8%

OpenStudy (anonymous):

compounded

OpenStudy (anonymous):

6% quarterly same yrs

OpenStudy (jim766):

A = 500,00( 1 + .08) ^(20/12) If you have a calculator that has parentheses , and can do Exponents...put it just like that If you are studying alg, a graphing calc would be good

OpenStudy (anonymous):

will look to see which i have

OpenStudy (anonymous):

buisness math

OpenStudy (anonymous):

just found one online lol

OpenStudy (anonymous):

thanks jim got answer